If the probability that the random variable X takes values x is given by P(X=x)=k(x+1)3−x,x=0,1,2,3,…, where k is a constant, then P(X≥2) is equal to :
Solutions · 1
Official solution
As, we know that sum of all the probabilities =1
\[
So, x=1∑∞P(X=x)=1⇒k[1+2⋅3−1+3⋅3−2+….∞]=1
\]
\[
Let S=1+32+323+….+∞⇒3S=0+31+322+333+….+∞
\]
On subtracting, we get
\[
32S=1+31+321+….+∞⇒32S=1−311=321⇒32S=23⇒S=49
\]
\[
So, k×49=1⇒k=94 Now, P(X≥2)=1−P(X<2)=1−P(X=0)−P(X=1)=1−94(1)−94×32=1−94−278=2727−12−8=277
\]